OpenStudy (anonymous):

One of the main contaminants of a nuclear accident, such as that at Chernobyl, is strontium-90, which decays exponentially at a rate of approximately 2.5% per year. (a) Write the ratio of strontium-90 remaining, p , as a function of years, t , since the nuclear accident.

OpenStudy (anonymous):

p(t)=?

OpenStudy (a_clan):

If p0 was the initial quantity, then quantity after time t, p(t) = p0 (e^ -kt) The ratio of strontium remaining,p= p(t)/p0 = (e^ -kt) = (e^ -0.025t)

OpenStudy (anonymous):

so it's just the last part? e^-....

OpenStudy (a_clan):

(e^ -0.025t) This is required function

OpenStudy (anonymous):

ok so p(t)= that

OpenStudy (a_clan):

No p(t) = p0 (e^ -kt) But the given problem is asking the ratio of Sr90 remaining, So, p= p(t)/p0 is the solution

OpenStudy (anonymous):

but it says p(t)=?

OpenStudy (a_clan):

That is just a difference of naming convention. If you want to call the ratio as p(t), then you can name 'the remaining quantity of Sr90' as , let us say 'pr'. Then, P(t) = pr/p0 = (e^ -0.025t)

OpenStudy (anonymous):

so its the e part i enter

OpenStudy (a_clan):

yes the entire e part (exponential part)

OpenStudy (anonymous):

okay thanks

OpenStudy (a_clan):

np

OpenStudy (anonymous):

Sorry, but i just entered e^....and it was wrong

OpenStudy (anonymous):

ugh

OpenStudy (a_clan):

There are different types of formula for the decay. y=a(1-r)^t is also used. try p(t)=(0.975)^t

OpenStudy (anonymous):

but it simply says p(t)= what? so I dont know what to enter.

OpenStudy (anonymous):

the e^... was wrong